OpenStudy (anonymous):

Find the exact value of:

OpenStudy (anonymous):

\[ \sin(\sin^{-1} (-1/2)+\cos^{-1} (-1/2))\]

OpenStudy (anonymous):

1 is the answer I got

OpenStudy (amistre64):

you mght have to use the sin formula for the addition of angles sin(a+b) = sina cosb + sinb cosa

OpenStudy (amistre64):

and a triangle labeled helps too

OpenStudy (anonymous):

first what we will do is that sin(-x) = -sinx and cos(-x) = cosx and also that cos theta =a and cos^-1 (a) = theta alright these are some rules

OpenStudy (anonymous):

@amistre64 , It is arcsin and not sin :D

OpenStudy (amistre64):

arctrigs are angles in disguise

OpenStudy (anonymous):

@amistre64 , my reply was in reference to this statement "you mght have to use the sin formula for the addition of angles sin(a+b) = sina cosb + sinb cosa"

OpenStudy (amistre64):

a = arcsin(-1/2) ; b = arccos(-1/2) they are still angles to work out

OpenStudy (amistre64):

sin(a+b) = sina cosb + sinb cosa = sin(arcsin(-1/2)) cos(arccos(-1/2)) + sin(arccos(-1/2)) cos(arcsin(-1/2))

OpenStudy (anonymous):

Why complicate things ? @amistre64 We should keep two things in mind 1) arcsin(x) always lies between (-pi/2 , pi/2) 2) arccos(x) always lies between (0 , pi) and just apply them

OpenStudy (anonymous):

I mean that they are the principal solutions ofcourse :D

OpenStudy (anonymous):

so in this case sin(sin−1(−1/2)+cos−1(−1/2)) so let us do this in steps and solve sin(sin−1(−1/2) first ok and arcsin(-1/2) = theta and sin theta = -1/2 and this is true when theta = 7pi/6 and now sin(7pi/6) = -1/2 and now we we will do cos−1(−1/2) = theta and cos theta = -1/2 and this is only true when theta = 2pi/3 and now let us add these 2 terms to get -1/2 +2pi/3 = (4pi -3)/6 :)

OpenStudy (anonymous):

for sin(sin−1(−1/2) we could have used a defintion where sin(arcsin(x)) this is just equal to x

OpenStudy (amistre64):

hamza; arc sin doesnt have 7pi/6 in its range ....

OpenStudy (anonymous):

why nott

OpenStudy (anonymous):

sin 7pi/6 = -1/2 ?

OpenStudy (anonymous):

This is probably the easiest, \(\arcsin x + \arccos x = \frac \pi 2 \) then the required exact value is 1.

OpenStudy (anonymous):

OMG lol i read the question totally wrong i thought it was sin(sin−1(−1/2) + cos-1(-1/2)

OpenStudy (anonymous):

i am so sorry

OpenStudy (anonymous):

NOTE: \(\large \arcsin( x) + \arccos( x) = \frac \pi 2 \forall x \in [-1,1] \)

OpenStudy (anonymous):

@Foolformath, I wanted to confirm my answer. Thanx to everyone for their input

OpenStudy (anonymous):

Glad to help :)